OpenStudy (yacoub1993):

@satellite73 Find the reference angle for the given angle. -404° A. 44° B. 46° C. 134° D. 136°

OpenStudy (rajat97):

yup i'm here

OpenStudy (yacoub1993):

HOW CAN I SOLVE THIS ONE?

OpenStudy (rajat97):

look if we see magnitude-wise, 404 is a greater number than 360 so the rotating arm of the angle will cover an angle more than 360 so just do it the reference angle = 404-360=44 deg

OpenStudy (yacoub1993):

it was -404 but your calculations are 404 - 360 = 44

OpenStudy (rajat97):

yeah we just want the reference angle so we take the magnitude i.e. the number that is present there and not the sign

OpenStudy (yacoub1993):

ohh ok

OpenStudy (yacoub1993):

i have another complicated question

OpenStudy (rajat97):

yeah but i cannot help you too long as i live in india and its 23:10 now

OpenStudy (yacoub1993):

OKAY

OpenStudy (yacoub1993):

THNXS... I APPRECIATED YOUR HELP THNXS

OpenStudy (rajat97):

yeah thanks for the medals that you have given me today but you can ask some questions now

OpenStudy (yacoub1993):

The wind chill factor represents the equivalent air temperature at a standard wind speed that would produce the same heat loss as the given temperature and wind speed. One formula for computing the equivalent temperature is

OpenStudy (rajat97):

my internet connection has some problem so please don't mind it if i go offline and come online

OpenStudy (yacoub1993):

ok

OpenStudy (rajat97):

i don't know what exactly s wind chill so i'm trying to work on it let's see what do i get

OpenStudy (rajat97):

i have no idea about it

OpenStudy (rajat97):

okay then let's meet tomorrow @17:00 hrs (indian standard time)